the total of three sister's age is 39. Dina is half as old as Michelle and three years younger than Juliette. How old are the sisters?
Answer by JulietG(1812)

D + M + J = 39
D =.5M, or M = 2D
D = J-3, or J = D+3
Take the known values from the bottom equation and put them into the top
D + M + J = 39
D + (2D) + (D+3) = 39
4D + 3 = 39
Subtract 3 from each side
4D = 36
Divide each side by 4
D = 9
If Dina is 9, then Michelle is 18, and Juliette is 12
9 + 18 + 12 = 39